OpenStudy (anonymous):

What is the solution set of the following system of equations? 5x - 3y - 10 = 0 x + y - 7 = 0 {(31/8, 25/8)} {(25/8, 31/8)} {(45/2, 31/2)} {(31/2, 45/2)}

OpenStudy (anonymous):

you have two unknowns and two equations rearrange one equation and plug it into the second

OpenStudy (anonymous):

for example x+y-7=0 so x= 7-y where you see x in equation one put in (7-y) and then solve for y once you have that find x

OpenStudy (anonymous):

y=25/8

OpenStudy (anonymous):

yes and use either equation to find x Good job

OpenStudy (anonymous):

31/8

OpenStudy (anonymous):

Solve this system algebraically. 7x - 2y = 4 5y + 3x = 10 {(-41/40, -41/58)} {(41/40, 41/58)} {(-40/41, -58/41)} {(40/41, 58/41)}

OpenStudy (anonymous):

same thing really, it's just a little tougher when you plug it in

OpenStudy (anonymous):

the only part i dont get is when you have to rearrange the problem

OpenStudy (anonymous):

it's an expansion of what you just did so let's rearrange the first one for y \[y= \frac{ 7x-4}{ 2 }\]

OpenStudy (anonymous):

x=20/38

OpenStudy (anonymous):

check it again

OpenStudy (anonymous):

10/19